OpenStudy (anonymous):

1. Find the slope for the values in the table below. x, -1, 0, 1, 2 y, 5, 3, 1, -1 A. 1 B. -1 C. 2 D. -2 2. What is the slope for the following table of values? The table represents the total amount spent in a paperback book club after an introductory fee of $12. # books purchased: 1/ 2/ 3 / 4 Cost of book club: 21/30/39/48 A. 3 B. 5 C. 7 D. 9

OpenStudy (fretje):

1/ slope is defined by the derived of the function or dy/dx = -2/1 = -2 (if you draw the line you can see that the y value decreases 2 each time the x value increases 1 ) so answer D) 2/ again the slope is the amount paid for one book : (48-12)/4 = 9 and (39-12)/3 = 9 and so on or you calculate the slope directly from its definition of the slope of a function slope = dy/dx = (48 - 21)/(4-1)=9 so answer D)